OpenStudy (anonymous):

3. What is the approximate volume of the sphere if the surface area is 482.8 mm²?

OpenStudy (cwrw238):

482 * (4/3) pi r^3 ---------- 4 pi r^2

OpenStudy (anonymous):

Elaborate.

OpenStudy (cwrw238):

= 482.8 * r / 3

OpenStudy (anonymous):

Take the SA equation solve for radius, plug radius into volume equation?

OpenStudy (anonymous):

I dont get it...

OpenStudy (cwrw238):

first find r from surface area equation = 4 pi r^2 then plug it into 482.8 * r / 3

OpenStudy (anonymous):

\[SA = 4\pi r ^{2}\] \[Volume= 4/3\pi r ^{3}\]

OpenStudy (anonymous):

I already know the surface area...its 482.8...i dont know the radius

OpenStudy (anonymous):

I dont know how to find the radius

OpenStudy (anonymous):

SA/4(PI) = r^2 take the sqrt(SA/4(PI) = r.

OpenStudy (anonymous):

thanks so much